1/2

Christian Morgenstern,  Christian Morgenstern,
The Banshee (An Approach)  Der Werwolf

tradukita de Max Knight

One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messingknauf und sprach zum Wolf, der seine Pfoten geduldig kreuzte vor dem Toten:

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

�Der Werwolf“ – sprach der gute Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s nennt, den Werwolf, - damit hat’s ein End!“

The banshee marveled at the cases and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

Dem Werwolf schmeichelten die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl noch!“

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

Der Dorfschullehrer aber musste

gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

…  …

2/2

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

http://www.poezio.net/version?poem-id=825&version-id=1660,1652